OpenStudy (anonymous):

Solve each inequality 8<2v-4<16

jimthompson5910 (jim_thompson5910):

the idea is to isolate v

jimthompson5910 (jim_thompson5910):

so we first add 4 to all sides then we divide all sides by 2

OpenStudy (anonymous):

Wait ill tey to answer

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

6<v<10?

jimthompson5910 (jim_thompson5910):

good, here are the steps if you're curious \[\large 8 < 2v-4 < 16\] \[\large 8+4 < 2v-4+4 < 16+4\] \[\large 12 < 2v+0 < 20\] \[\large 12 < 2v < 20\] \[\large \frac{12}{2} < \frac{2v}{2} < \frac{20}{2}\] \[\large 6 < \frac{1v}{1} < 10\] \[\large 6 < 1v < 10\] \[\large 6 < v < 10\]

jimthompson5910 (jim_thompson5910):

I'm sure you did all that

OpenStudy (anonymous):

Yeah im right. But i did not that 1 i just cancel it out

jimthompson5910 (jim_thompson5910):

that's ok, I put more steps than needed really to be really technical

OpenStudy (anonymous):

-7d <_4d-3<_-1

jimthompson5910 (jim_thompson5910):

hint: use the idea that x < y < z breaks up into x < y and y < z

OpenStudy (anonymous):

How can i use that in this equation whats the y?

jimthompson5910 (jim_thompson5910):

-7d <_4d-3<_-1 breaks up into -7d <_4d-3 and 4d-3<_-1

jimthompson5910 (jim_thompson5910):

solve each for d

OpenStudy (anonymous):

I dnt know how to solve this one its complicated

jimthompson5910 (jim_thompson5910):

give it a shot

jimthompson5910 (jim_thompson5910):

tackle each one on its own

OpenStudy (anonymous):

Ill try

OpenStudy (anonymous):

d>_d<_-4?

jimthompson5910 (jim_thompson5910):

hmm that seems off

OpenStudy (anonymous):

Dnt know pls explain thnx :)

jimthompson5910 (jim_thompson5910):

what do you get when you solve -7d <_4d-3

OpenStudy (anonymous):

-d<d

jimthompson5910 (jim_thompson5910):

hint: -7d <_4d-3 -7d -4d <_ -3

OpenStudy (anonymous):

d<_-3?

jimthompson5910 (jim_thompson5910):

you're missing something on the left side

OpenStudy (anonymous):

-3?

jimthompson5910 (jim_thompson5910):

no -7d - 4d = ??

OpenStudy (anonymous):

-11

OpenStudy (anonymous):

How about the other side?

jimthompson5910 (jim_thompson5910):

so we have -11d <_ -3 now divide both sides by -11 to isolate d, this will flip the inequality sign

OpenStudy (anonymous):

<_ 3.66

OpenStudy (anonymous):

>_ imean

jimthompson5910 (jim_thompson5910):

leave it as a fraction

OpenStudy (anonymous):

I got 3.6

jimthompson5910 (jim_thompson5910):

leave it as a fraction to get d >= 3/11

jimthompson5910 (jim_thompson5910):

now solve the other piece

OpenStudy (anonymous):

Ow i got >_ 2d

jimthompson5910 (jim_thompson5910):

solve 4d-3<_-1 for d

OpenStudy (anonymous):

Yeah i put 3 in both sides

OpenStudy (anonymous):

Then the answer was divided into 4 so i got 2

jimthompson5910 (jim_thompson5910):

2/4 is not 2 though

OpenStudy (anonymous):

Why

jimthompson5910 (jim_thompson5910):

4/2 = 2 not the other way around

jimthompson5910 (jim_thompson5910):

4/2 = 2 2/4 = ???

OpenStudy (anonymous):

.5

jimthompson5910 (jim_thompson5910):

or 1/2

jimthompson5910 (jim_thompson5910):

so d <= 1/2

jimthompson5910 (jim_thompson5910):

so all together, we know this d >= 3/11 and d <= 1/2 put together, the final answer is \[\large \frac{3}{11} \le d \le \frac{1}{2}\]

OpenStudy (anonymous):

So ill combine and thats the final asnwer

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

This is the equation i cant understand

jimthompson5910 (jim_thompson5910):

ok last one

OpenStudy (anonymous):

6y-3<27 or -4y+2<-26 what ir means

jimthompson5910 (jim_thompson5910):

you solve each inequality for y so what do you get when you solve 6y-3<27

OpenStudy (anonymous):

Hold on a sec

OpenStudy (anonymous):

Y<5

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

solve -4y+2<-26 for y

OpenStudy (anonymous):

Y<-6

jimthompson5910 (jim_thompson5910):

no that's incorrect

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

-7?

OpenStudy (anonymous):

Is that right?

jimthompson5910 (jim_thompson5910):

it should be y > 7

jimthompson5910 (jim_thompson5910):

so after solving 6y-3<27 or -4y+2<-26 , we get this y < 5 or y > 7

OpenStudy (anonymous):

So thats the final answer?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Thank you so much

jimthompson5910 (jim_thompson5910):

you're welcome